Saturday, September 16, 2006

Timing delay for target nozzle

In this week, we tried to focus the laser beam to the Xe cluster for measuring the Xe(L) emission. Because the nozzle has been changed, the timing delay between laser pulse and nozzle trigger must be adjusted. On Monday this delay time was about 478 us, however it changed to 410 us next day, on Friday the dealy was 491us. The range is around 10 to 80 us, beyand these values, we saw nor Xe(M) light on the top CCD neither any singals from crystal spectrometer (for measuring L-shell light).

No comments: